OpenStudy (anonymous):

integral of 0 to 1 integral of sqrt(x) to 1 of sin(y^3) dy dx (double integral problem... please help! im not getting anywhere!!!!)

OpenStudy (xishem):

\[\int\limits_{0}^{1} \int\limits_{\sqrt{x}}^{1}\sin(y^3)\ dy\ dx\]This is a difficult integral. That's usually a sign that you need to try changing your order of integration. Draw a sketch, change your limits, and try integrating w.r.t. x first.